HOME ESR EXPERIMENTS LABS etc
E121 E127 E132 E125 E143 laser_cooling_2021 E142 Ê128 E146 E0052 E0018 E0028 E0038 E0075 E0203_E0018
  masses & isomeres 2025, Page 3 of 7  Not logged in ELOG logo
ID Date Author Category Subject Text Attachments
  95   Wed Apr 2 22:47:25 2025 ID, HA (Radio Shahab as background)GeneralLassie and Schottky TV screentshotsWe managed to take some screenshots while
Shahab harrassed us with AI-created "RammstAIn"
music... happy to end my shift now!
 2025-04-02_22-44-18-420.png2025-04-02_22-45-12-849.png2025-04-02_22-55-14-900.png 
  94   Wed Apr 2 21:22:41 2025 ID, HAGeneralLassie Come-Home Screenshots (wuff!)(Who let the dog out?)
Sharing with FRS 6:1 since a while now.
 2025-04-02_21-25-00-716.png2025-04-02_21-25-42-535.png2025-04-02_21-26-12-022.png 
  93   Wed Apr 2 20:57:34 2025 ShahabDetectors10GB cables for the DAQ rackWe have been working together with IT on the
speed of data transfer: with the new 10 GB
interface we can now transfer data from NAS
  
  92   Wed Apr 2 20:15:29 2025 ID, HAGeneralLassie screenshots + Schottky TVScreeeeeeeenshooooots!  2025-04-02_20-15-48-827_tcl1064.png2025-04-02_20-17-14-017.png2025-04-02_20-18-29-150.png 
  91   Wed Apr 2 19:13:50 2025 ID, HAGeneralLassie screenshots and Schottky screenshotLooks a bit more regular now (FRS people having
dinner???)
 2025-04-02_19-10-29-388_tcl1064.png2025-04-02_19-14-53-428_tcl1146.png2025-04-02_19-19-23-765.png 
  90   Wed Apr 2 18:31:52 2025 IDGeneralFYI for morning shifts ThuThe PLEIADES team (pocket detectors) has discussed
with Wolfram to get access to the ESR tomorrow
(Thu) morning ~8:00 to check their detectors
  
  89   Wed Apr 2 18:22:30 2025 IDGeneralSpill cycle distributionThe spill cycle distributions is absolutely
nuts since beginning of the shift. On the
positive side: we got a lot of spills since
  
  88   Wed Apr 2 18:15:06 2025 ID, HA, MWGeneralLassie screenshots   2025-04-02_18-14-21-624.png 
  87   Wed Apr 2 18:13:20 2025 ID, HAGeneralRSA spectra   IMG_0087.jpeg 
  86   Wed Apr 2 17:46:12 2025 Ruijiu, Wolfram, Xinliang, WangqianAnalysisThe spectrum of 194Pb82+ vs 184Pt78+In summary, the statistical results collected
between 5:00 and 12:00 AM on April 2, 2025.
For each injection, we calculate the frequency
 photo_2025-04-02_17-47-24.jpgplot_fft.cxxScreenshot_2025-04-02_22-00-57_1.png 
  85   Wed Apr 2 17:06:45 2025 ID, HA et alGeneralLassie screenshots   2025-04-02_17-09-17-201.png 
  84   Wed Apr 2 16:28:14 2025 SS, ID, HA etcDAQNTCAP stop/startNTCAP was stopped and restarted.   
  83   Wed Apr 2 15:46:24 2025 ID, HA,MWGeneralBeam back in ESR, beam is back, we started DAQ again. All files
created after 15:45 should be good from now
on.
 20250402_154135.jpg 
  82   Wed Apr 2 15:07:52 2025 Michael, Alex, MaeveAcceleratorStill not a good beam in the ESRThe (power supply of) SIS kicker magnet has
been repaired and the beam has been restored
from about 14:20.
  
  81   Wed Apr 2 13:35:17 2025 Michael, Alex, MaeveDAQRSAs data takingSince there will be no beam for about one
hour, the data taking in RSAs has been stopped.
  
  80   Wed Apr 2 13:27:52 2025 Michael, Alex, MaeveAcceleratorNo beam in the ESRThere is no beam in the ESR since couple of
minutes. According to the info from HKR,
the experts started to repair the magnets
  
  79   Wed Apr 2 12:58:15 2025 Ruijiu, WolframAnalysisfrequency drift   photo_2025-04-02_15-24-34.jpgphoto_2025-04-02_15-28-44.jpgfrequency_194Pb82.xlsx 
  78   Wed Apr 2 12:13:57 2025 Michael, AlexAcceleratorSome empty injections triggered while beam patterns are changedOperators seem to play with spill cycles,
we get every 1:1 shot, then 1:10 again. During
this, we noticed a few "empty injections"
  
  77   Wed Apr 2 11:09:42 2025 Michael, AlexAcceleratorThe beam is back in the ESRWe have the beam again the ESR since around
11:00.
The Schottky spectra look the same as before.
  
  76   Wed Apr 2 10:55:16 2025 Michael, Alex, WolframAcceleratorNo beam in the ESRSince few minutes do not see any beam in the
ESR (on RSAs). There is beam in SIS and injection
requests from the ESR seem to be there as
  
ELOG V3.1.5-fc6679b